# A Modified 3D Biconical Outflow Model: Spatial Constraints on AGN-driven Outflows

### v1

@@ 4.2 Constraints on $V_{max}$ and $\theta_{out}$ Distributions
Figure 5: Left: Kullback-Leibler divergence values depending on the PDF of the launching velocity. The PDF of the opening angle is fixed to the best. The mean and standard deviation of $V_{max}$ first determine the Gaussian distribution, and we use the truncated Gaussian from 100 km s-1as a PDF of the $V_{max}$. The minimum Kullback-Leibler divergence is found when the mean and standard deviation are 250 and 400 km s-1. Right: The same plot but $D_{max}$ from the 2D KS test. The minimum value is found when the mean and standard deviation are 350 and 300 km s-1.
@@
Figure 6: Left: Kullback-Leibler divergence values depending on the PDF of the opening angle. The minimum Kullback-Leibler divergence is found when the mean and standard deviation are 30 and 10, respectively. Right: The same plot but $D_{max}$ from the 2D KS test. The minimum value is found when the mean and standard deviation are 30 and 5.
@@
We first investigate variations of Kullback-Leibler divergence and $D_{max}$ values depending on the input launching velocity and outer opening angle PDFs in Figure 5 and 6. In each map, we fix the PDF of the other parameter to the best one. The Kullback-Leibler divergence and $D_{max}$ maps of the launching velocity show local minima near the center, ruling out the lower left and upper right corners. This indicates that the observed VVD distribution prefers a moderate range of launching velocity in general. On the other hand, the PDF of the opening angle prefers small values (i.e., lower left corner regions in Figure 6). Specifically, Kullback-Leibler divergence is minimized when the launching velocity follows ($\mu_{V},\sigma_{V}$)=(250 km s-1, 400 km s-1) and the opening angle follows ($\mu_{\theta},\sigma_{\theta}$)=(30, 10). $D_{max}$ exhibits a minimum value when the launching velocity follows ($\mu_{V},\sigma_{V}$)=(350 km s-1, 300 km s-1) and the opening angle follows ($\mu_{\theta},\sigma_{\theta}$)=(30, 5).
@@
We present the best mock VVD distributions based on Kullback-Leibler divergence (left) and $D_{max}$ (right) in Figure 7. The two best mock VVD distributions are generally consistent with the observed one, reproducing both the V-shape feature and the overall range of the V[OIII] and $\sigma_{\mathrm{[OIII]}}$. The best mock VVD distribution based on Kullback-Leibler divergence exhibits a slightly larger $\sigma_{\mathrm{[OIII]}}$ range than $D_{max}$. Despite the extinction limit ($A\geq 0.2$), the V-shape envelope of the mock VVD distribution is slightly broader than the observed one, especially in the low $\sigma_{\mathrm{[OIII]}}$ regime. This may result from several factors, such as large observational uncertainties for weak outflow AGNs.
@@
Figure 7: The best mock VVD distributions based on the Kullback-Leibler divergence (left) and $D_{max}$ (right). The contours from the observed VVD distribution are overlaid. Dashed box indicates a selection criterion of strong outflow AGNs ($|\mathrm{V}_{\mathrm{[OIII]}}|>300$ km s-1or $\sigma_{\mathrm{[OIII]}}>400$ km s-1).
@@
Figure 8: The best PDFs of the launching velocity based on the Kullback-Leibler divergence (left) and $D_{max}$ (right). The histograms show the best PDFs for the entire mock VVD distributions, which are derived from Gaussian distributions truncated from 100 km s-1. The mean and standard deviation of the Gaussian are denoted in each panel.
@@
Figure 9: The launching velocity distributions for a subset of strong outflow targets (outside the dashed boxes in Figure 7). The distributions are obtained from the best mock VVD distributions based on Kullback-Leibler divergence (left) and $D_{max}$ (right).
@@
Figure 10: The best PDFs of the opening angle based on the Kullback-Leibler divergence (left) and $D_{max}$ (right). Both follow Gaussian distributions truncated from 30 to 70.
@@
The histograms in Figure 8 show the best PDFs of the launching velocity based on Kullback-Leibler divergence (left) and $D_{max}$ (right). The two quantities adopt similar best PDFs for the launching velocity. As predicted in Section 4.1, the launching velocities are mostly smaller than 1000 km s-1, while 2 - 5% exhibit higher launching velocities (1000 - 1500 km s-1).
@@
Lastly, the best PDFs of the opening angle (Figure 10) are slightly different depending on the statistical quantities. Specifically, $\theta_{out}\geq 50^{\circ}$ is only exhibited in the best PDF based on Kullback-Leibler divergence, while the best PDF based on $D_{max}$ is concentrated on $\theta_{out}=30^{\circ}-40^{\circ}$. Nevertheless, the two best PDFs consistently favor relatively narrow outer opening angles, indicating that the opening angle of the local type 2 AGN is likely to be narrow, mostly 30 - 40.
4.3 Testing Outflow Sizes
@@
Figure 11: An example of mock IFU analysis when redshift is 0.05. Left: Radial profiles of $\sigma_{\mathrm{[OIII]}}$ (yellow markers) and W80 (blue markers). Large colored markers show the means of velocity widths in individual radial bins. Blue and yellow horizontal lines represent the criteria of the outflow region for W80 (600 km s-1) and $\sigma_{\mathrm{[OIII]}}$ ($\sigma_{*}$), respectively. Vertical dashed lines indicate half-width-half-maximum (HWHM) seeing size (red), input outflow size (black), the W80-based size (blue), and the $\sigma_{\mathrm{[OIII]}}$-based size (yellow). Right: 2D maps of velocity dispersion and W80. The colors of dashed line circles follow the same color scheme as vertical lines in the left panel.
Figure 12: The same plots as Figure 11 but when redshift is 0.3.
@@
We present the spatial distributions of velocity widths of the two examples in Figure 11 and 12. In our previous IFU studies, radial profiles of velocity dispersion often exhibit constant values in the central region (i.e., initial plateau) before a smooth decrease (e.g., Karouzos16a; Kang18). This initial plateau is well reproduced in our mock IFU data, for both $\sigma_{\mathrm{[OIII]}}$ (yellow) and W80 (blue). Moreover, the radius of the initial plateau is comparable to the HWHM of the seeing (red dashed line), indicating that it mainly results from the seeing effect.
@@
5 Discussion
5.1 Comparison with Other Outflow Models
@@
In this section, we discuss the characteristics of our model in comparison with the previous models developed over the last two decades. First, early outflow models did not include a rotating disk. For example, Crenshaw00a; Crenshaw00b constructed a hollow bicone model to reproduce position-velocity (PV) diagrams of nearby AGNs (e.g., Das05; Das06). Subsequently, Crenshaw10a; Crenshaw10b added a dust extinction component to the model to apply for dozens of nearby AGNs (e.g., Fischer10; Fischer11; Fischer13; Meena21; Falcone24). Since these models were developed for the HST observations of nearby AGNs, it is not essential to include the effect of the point spread function (PSF).
@@
On the other hand, Carniani15 developed a model with a single approaching cone for testing their spectroastrometric analysis of powerful outflows in quasars at cosmic noon. Since the seeing effect is substantial for their high-z AGNs, they convolved each velocity channel map with the PSF, reproducing consistent [O III] flux, velocity, and velocity dispersion maps with the observations. Although their model lacks a disk component, the contribution of the disk component is presumably weak in these extremely luminous AGNs.
@@
Second, other models were also reported for accommodating various kinematical considerations. For example, Muller-Sanchez11 constructed a model to analyze the coronal line regions (CLRs) using AO-assisted near-infrared IFU data. By combining a bicone model with a rotating disk, they successfully fitted CLR kinematics at the central several hundred pc, while no seeing effect was included.
@@
Recently, Marconcini23 developed a new model (MOKA3D) based on Venturi17 by constructing a 3D distribution of point-like clouds to reproduce the observed IFU data. The model also convolves the 3D data and assigns a proper weight to each cloud. Since outflows often exhibit clumpy regions within their bicone geometry (e.g., Kakkad23), their model has a major advantage in reproducing various substructures, showing excellent agreements with the observations of both nearby and high-z AGNs (Cresci23; Marconcini23; Marconcini25a; Marconcini25b).
@@
Compared to the previous models, which are mainly focused on fitting 3D data cubes of individual AGNs, our model is better suited for investigating the distributions of outflow parameters for a large sample, including the effect of seeing size and the disk component. Note that our model can also be utilized for reproducing the observation of individual AGNs (e.g., Shin19). Thus, our analysis with the modified model is complementary to the previous outflow models.
5.2 Constraints on the AGN-driven outflows
@@
The launching velocity is the key parameter for quantifying outflow energetics (see Section 5.4), while the opening angle is related to how strongly gas outflows interact with the ISM. Thus, it is essential to constrain the launching velocity and opening angle of outflows for a large sample to investigate AGN feedback.
@@
We report launching velocities of $\lesssim$ 1500 km s-1  and typical opening angles of 30-40 (Section 4.2). For a consistency check, we discuss the previous estimations of these two parameters, mostly from nearby AGNs. For example, Crenshaw00a; Crenshaw00b reported the inner/outer opening angles as 26/40 and the maximum outflow velocity as 1300 km s-1for NGC 1068, and inner/outer opening angles as 20/36 and the maximum outflow velocity as 750 km s-1for NGC 4151 (see also Das05; Das06). After including dust extinction, the outer opening angles are slightly changed to 40 for NGC 1068 and 33 for NGC 4151 (Crenshaw10b).
@@
Their following studies also reported outflow parameters of other nearby AGNs (Crenshaw10a; Fischer10; Fischer11; Meena21; Falcone24). In particular, among the cases of 17 Seyfert galaxies investigated by Fischer13, the maximum velocity agrees with our range of $V_{max}\leq 1500$km s-1, except for NGC 1068. On the other hand, the outer opening angle distribution is slightly different. While a majority of their sample shows consistent opening angles (<<50) with our results, 7 AGNs have wider opening angles. The difference can be caused by a couple of effects. First, AGNs with higher luminosity and stronger outflows are likely to have wider opening angles (Bae16). For example, only 13% of our type 2 AGN sample have bolometric luminosity higher than $10^{44.5}$ erg/s, while more than half of their sample (9/17) are more luminous than $10^{44.5}$ erg/s. Second, their sample was selected based on clear signs of extended outflows, whereas we do not apply any criterion on the outflow extent. The inclusion of the disk component in our model can also induce differences.
@@
The maximum velocity and the outer opening angle estimated by Muller-Sanchez11; Muller-Sanchez16 are also consistent with our constraints. On the other hand, for the best-fit parameters from Nevin18, while the maximum velocities ($V_{max}=230-720$ km s-1) also agree with our study, most of the opening angle values are much higher than the typical range of our constraint. This discrepancy also appears to be due to different biconical model and/or the characteristics of their sample (i.e., AGNs with double-peaked narrow emission lines). Other studies that constrained the opening angles of individual galaxies (e.g., Baron18; Riffel21; Costa-Souza24) are in agreement with our constraint.
5.3 True Outflow Size
@@
In Section 4.3, we verified that the outflow size measured based on velocity widths can be significantly overestimated by the seeing, especially when a true outflow size is comparable to or smaller than the seeing size. To properly constrain the spatial scale of outflows, it is essential to investigate the seeing-limited cases in detail.
Previous studies reported notable spatial features in luminous type 2 AGNs, i.e., circularly extended NLR morphologies or enhanced W80 values with flat radial profiles (e.g., Liu13a; Liu13b; Harrison14). Interestingly, our model reproduces a similar flat W80 profile when the outflow is compact and the disk emission is much weaker than the outflow. Seeing convolution of a compact outflow generates a broad wing component with an almost constant velocity width across the entire FoV, while the disk emission is too faint to affect the velocity width, even at pixels far from the center. Thus, given that their samples consist of luminous type 2 quasars, implying relatively weak disk emission, the flat W80 profiles may indicate smaller outflow sizes.
Furthermore, Villar-Martin16 found that only 2 out of 15 AGNs in their sample show clearly resolved outflows by comparing the spatial profiles of seeing with the [O III] wing emission. Based on spectroastrometry and seeing sizes, they constrained lower and upper limits of the outflow sizes as several hundred pc and a few kpc, respectively. These results are in agreement with our constraints and represent observational evidence of the seeing effect in local AGNs.
@@
In our previous IFU studies (Karouzos16a; Bae17; Kang18; Luo21; Kim23), the outflow sizes were determined based on the $\sigma_{\mathrm{[OIII]}}$-based method. Our main conclusion was that the outflow sizes are mostly less than several kpc scales, indicating a lack of galactic scale feedback. We also reported a kinematic outflow size–luminosity relation between the outflow size and the [O III] luminosity with a slope of $\sim$0.3. For this work, we subtracted the HWHM of the seeing size from the outflow size in quadrature to correct for the seeing effect. However, this correction may not be sufficient as shown in Section 4.3. Thus, our measurements can be conservatively considered as upper limits, particularly for higher redshift targets among our sample ($z>0.1$).
To test the reliability of the size measurements, we apply the $\sigma_{\mathrm{[OIII]}}$-based method to various mock data. For this test, we set the redshift z=0.1 as a representative case and restrict $V_{max}$ from 750 to 2000 km s-1 to focus on strong outflows (see Section 4.2). In Figure 13, we present the measured outflow size as a function of the input (true) size. As expected, the outflow size is overestimated by more than a factor of two when the true outflow size is smaller than the FWHM of the seeing ($\sim$1.52 kpc at z = 0.1). On the other hand, when the seeing is smaller than the true size, the measured size is generally consistent with the input size with slight underestimations (<< 0.3 dex). Therefore, the $\sigma_{\mathrm{[OIII]}}$-based method provides a good size estimation within 0.3 dex, or at least can be used as a reliable upper limit, depending on the relative seeing size.
We emphasize that although the $\sigma_{\mathrm{[OIII]}}$-based method overestimates the sizes of compact outflows, the main conclusion of Kim23, i.e., lack of global feedback, remains valid. Since the true size can be more compact than the measured one, the upper limit of the size still provides strong constraints on the spatial scales of AGN feedback.

### v2

@@ IV.2 Constraints on $V_{max}$ and $\theta_{out}$ Distributions
Figure 5: Left: KL divergence values for differently assumed PDFs of the launching velocity $V_{\rm max}$, while keeping the opening angle distribution fixed to its best-fit value. The Gaussian parameters (mean and standard deviation) for $V_{\rm max}$ are varied and truncated below 100 km s-1. The minimum KL divergence is found at a mean of 250 km s-1 and a standard deviation of 400 km s-1. Right: same plot but showing $D_{\rm max}$ values instead of KL divergence. In this case, the minimum occurs at a mean of 350 km s-1 and a standard deviation of 300 km s-1. In each panel, the location of the lowest value is highlighted by a dashed-line box.
@@
Figure 6: Left: KL divergence values computed for differently assumed PDFs of the opening angle $\theta_{\rm out}$. The minimum KL divergence is found when the mean and standard deviation are $30^{\circ}$ and $10^{\circ}$, respectively. Right: Same as left, but showing the $D_{\rm max}$ values from the 2D KS test. The minimum occurs at a mean of $30^{\circ}$ and a standard deviation of $5^{\circ}$. In each panel, the location of the lowest value is highlighted by a dashed-line box.
@@
We first investigate variations in the KL divergence and $D_{max}$ values depending on the input launching velocity and outer opening angle PDFs in Figures 5 and 6. For visualization, we collapse the 4D arrays into 2D maps by fixing the other parameter to its best PDF. The KL divergence and $D_{max}$ maps of the launching velocity show local minima near the center, ruling out the lower left and upper right corners. This indicates that the observed VVD distribution generally prefers a moderate range of launching velocity. On the other hand, the opening angle PDF favors small values (i.e., lower left corner regions in Figure 6).
@@
We present the best mock VVD distributions based on KL divergence (left) and $D_{max}$ (right) in Figure 7. The two best mock VVD distributions are generally consistent with the observed one, reproducing both the V-shape feature and the overall range of the V[OIII] and $\sigma_{\mathrm{[OIII]}}$. The best mock VVD distribution based on KL divergence exhibits a slightly larger $\sigma_{\mathrm{[OIII]}}$ range than $D_{max}$. Despite the extinction limit ($A\geq 0.2$), the V-shape envelope of the mock VVD distribution is slightly broader than the observed one, especially in the low $\sigma_{\mathrm{[OIII]}}$ regime. This may result from several factors, such as large observational uncertainties for weak outflow AGNs.
@@
Figure 7: Best-match mock VVD distributions selected based on KL divergence (left) and $D_{\rm max}$ (right). The contours from the observed VVD distribution are overlaid for comparison. The dashed-box outlines the selection criterion for “strong outflow” AGNs, defined as $|\mathrm{V}_{\rm[OIII]}|>300\ \mathrm{km\ s^{-1}}$ or $\sigma_{\rm[OIII]}>400\ \mathrm{km\ s^{-1}}$.
@@
Figure 8: Best-fitting PDFs of the launching velocity based on the KL divergence (left) and $D_{max}$ (right). Histograms show the adopted Gaussian PDFs for the full mock VVD distributions with truncation at 100 km s-1. The mean and standard deviation of each Gaussian are indicated in their respective panels.
@@
Figure 9: Distributions of the launching velocity for a subsample of strong-outflow AGNs (those lying outside the dashed boxes in Figure 7). Histograms are derived from the best-matching mock VVD distributions based on KL divergence (left) and $D_{\rm max}$ (right).
@@
Figure 10: Best-fit PDFs for the outflow opening angle ($\theta_{\rm out}$) based on the KL divergence (left) and $D_{max}$ (right). All distributions are truncated Gaussians defined over the range $30^{\circ}$$70^{\circ}$.
@@
The histograms in Figure 8 show the best PDFs of the launching velocity based on the KL divergence (left) and $D_{max}$ (right). The two quantities adopt similar best PDFs for the launching velocity. As predicted in Section IV.1, the launching velocities are mostly smaller than 1000 km s-1, while 2 - 5% exhibit higher launching velocities (1000 - 1500 km s-1).
@@
Lastly, the best PDFs of the opening angle (Figure 10) are slightly different depending on the statistical quantities. Specifically, $\theta_{out}\geq 50^{\circ}$ is only exhibited in the best PDF based on the KL divergence, while the best PDF based on $D_{max}$ is concentrated on $\theta_{out}=30^{\circ}-40^{\circ}$. Nevertheless, the two best PDFs consistently favor relatively narrow outer opening angles, indicating that the opening angle of the local type 2 AGN is likely to be narrow, mostly 30 - 40.
IV.3 Testing Outflow Sizes
@@
[…]
Figure 12: Same as Figure 11 but for a higher redshift of $z=0.3$.
@@
We present the spatial distributions of velocity widths of the two examples in Figure 11 and 12. In our previous IFU studies, radial profiles of velocity dispersion often exhibit constant values in the central region (i.e., initial plateau) before a smooth decrease (e.g., Karouzos et al., 2016; Kang and Woo, 2018). This initial plateau is well reproduced in our mock IFU data, for both $\sigma_{\mathrm{[OIII]}}$ (yellow) and W80 (blue). Moreover, the radius of the initial plateau is comparable to the HWHM of the seeing (red dashed line), indicating that it mainly results from the seeing effect.
@@
V Discussion
V.1 Comparison with Other Outflow Models
@@
In this section, we discuss the characteristics of our model in comparison with the previous models developed over the last two decades. First, early outflow models did not include a rotating disk. For example, Crenshaw and Kraemer (2000); Crenshaw et al. (2000) constructed a hollow bicone model to reproduce position-velocity (PV) diagrams of nearby AGNs (e.g., Das et al., 2005, 2006). Subsequently, Crenshaw et al. (2010a, b) added a dust extinction component to the model to apply for dozens of nearby AGNs (e.g., Fischer et al., 2010, 2011, 2013; Meena et al., 2021; Falcone et al., 2024). Since these models were developed for the HST observations of nearby AGNs, it is not essential to include the effect of the point spread function (PSF).
@@
On the other hand, Carniani et al. (2015) developed a model with a single approaching cone for testing their spectroastrometric analysis of powerful outflows in quasars at cosmic noon. Since the seeing effect is substantial for their high-z AGNs, they convolved each velocity channel map with the PSF, reproducing consistent [O III] flux, velocity, and velocity dispersion maps with the observations. Although their model lacks a disk component, the contribution of the disk component is presumably weak in these extremely luminous AGNs.
@@
Second, other models were also reported for accommodating various kinematical considerations. For example, Müller-Sánchez et al. (2011) constructed a model to analyze the coronal line regions (CLRs) using AO-assisted near-infrared IFU data. By combining a bicone model with a rotating disk, they successfully fitted CLR kinematics at the central several hundred pc, while no seeing effect was included.
@@
Recently, Marconcini et al. (2023) developed a new model (MOKA3D) based on Venturi et al. (2017) by constructing a 3D distribution of point-like clouds to reproduce the observed IFU data. The model also convolves the 3D data and assigns a proper weight to each cloud. Since outflows often exhibit clumpy regions within their bicone geometry (e.g., Kakkad et al., 2023), their model has a major advantage in reproducing various substructures, showing excellent agreements with the observations of both nearby and high-z AGNs (Cresci et al., 2023; Marconcini et al., 2023, 2025a, 2025b).
@@
Compared to the previous models, which are mainly focused on fitting 3D data cubes of individual AGNs, our model is better suited for investigating the distributions of outflow parameters for a large sample, including the effect of seeing size and the disk component. Note that our model can also be utilized for reproducing the observation of individual AGNs (e.g., Shin et al., 2019). Thus, our analysis with the modified model is complementary to the previous outflow models.
V.2 Constraints on the AGN-driven outflows
@@
The launching velocity is the key parameter for quantifying outflow energetics (see Section V.4), while the opening angle is related to how strongly gas outflows interact with the ISM. Thus, it is essential to constrain the launching velocity and opening angle of outflows for a large sample to investigate AGN feedback.
@@
We report launching velocities of $\lesssim$ 1500 km s-1 and typical opening angles of 30-40 (Section IV.2). For a consistency check, we discuss the previous estimations of these two parameters, mostly from nearby AGNs. For example, Crenshaw and Kraemer (2000); Crenshaw et al. (2000) reported the inner/outer opening angles as 26/40 and the maximum outflow velocity as 1300 km s-1 for NGC 1068, and inner/outer opening angles as 20/36 and the maximum outflow velocity as 750 km s-1 for NGC 4151 (see also Das et al., 2005, 2006). After including dust extinction, the outer opening angles are slightly changed to 40 for NGC 1068 and 33 for NGC 4151 (Crenshaw et al., 2010b).
@@
Their following studies also reported outflow parameters of other nearby AGNs (Crenshaw et al., 2010a; Fischer et al., 2010, 2011; Meena et al., 2021; Falcone et al., 2024). In particular, among the cases of 17 Seyfert galaxies investigated by Fischer et al. (2013), the maximum velocity agrees with our range of $V_{max}\leq 1500$km s-1, except for NGC 1068. On the other hand, the outer opening angle distribution is slightly different. While a majority of their sample shows consistent opening angles (<<50) with our results, 7 AGNs have wider opening angles. The difference can be caused by a couple of effects. First, AGNs with higher luminosity and stronger outflows are likely to have wider opening angles (Bae and Woo, 2016). For example, only 13% of our type 2 AGN sample have bolometric luminosity higher than $10^{44.5}$ erg/s, while more than half of their sample (9/17) are more luminous than $10^{44.5}$ erg/s. Second, their sample was selected based on clear signs of extended outflows, whereas we do not apply any criterion on the outflow extent. The inclusion of the disk component in our model can also induce differences.
@@
The maximum velocity and the outer opening angle estimated by Müller-Sánchez et al. (2011, 2016) are also consistent with our constraints. On the other hand, for the best-fit parameters from Nevin et al. (2018), while the maximum velocities ($V_{max}=230-720$ km s-1) also agree with our study, most of the opening angle values are much higher than the typical range of our constraint. This discrepancy also appears to be due to different biconical model and/or the characteristics of their sample (i.e., AGNs with double-peaked narrow emission lines). Other studies that constrained the opening angles of individual galaxies (e.g., Baron et al., 2018; Riffel, 2021; Costa-Souza et al., 2024) are in agreement with our constraint.
V.3 True Outflow Size
@@
In Section IV.3, we verified that the outflow size measured based on velocity widths can be significantly overestimated by the seeing, especially when a true outflow size is comparable to or smaller than the seeing size. To properly constrain the spatial scale of outflows, it is essential to investigate the seeing-limited cases in detail.
Previous studies reported notable spatial features in luminous type 2 AGNs, i.e., circularly extended NLR morphologies or enhanced W80 values with flat radial profiles (e.g., Liu et al., 2013a, b; Harrison et al., 2014). Interestingly, our model reproduces a similar flat W80 profile when the outflow is compact and the disk emission is much weaker than the outflow. Seeing convolution of a compact outflow generates a broad wing component with an almost constant velocity width across the entire FoV, while the disk emission is too faint to affect the velocity width, even at pixels far from the center. Thus, given that their samples consist of luminous type 2 quasars, implying relatively weak disk emission, the flat W80 profiles may indicate smaller outflow sizes.
Furthermore, Villar-Martín et al. (2016) found that only 2 out of 15 AGNs in their sample show clearly resolved outflows by comparing the spatial profiles of seeing with the [O III] wing emission. Based on spectroastrometry and seeing sizes, they constrained lower and upper limits of the outflow sizes as several hundred pc and a few kpc, respectively. These results are in agreement with our constraints and represent observational evidence of the seeing effect in local AGNs.
@@
In our previous IFU studies (Karouzos et al., 2016; Bae et al., 2017; Kang and Woo, 2018; Luo et al., 2021; Kim et al., 2023), the outflow sizes were determined based on the $\sigma_{\mathrm{[OIII]}}$-based method. Our main conclusion was that the outflow sizes are mostly less than several kpc scales, indicating a lack of galactic scale feedback. We also reported a kinematic outflow size–luminosity relation between the outflow size and the [O III] luminosity with a slope of $\sim$0.3. For this work, we subtracted the HWHM of the seeing size from the outflow size in quadrature to correct for the seeing effect. However, this correction may not be sufficient as shown in Section IV.3. Thus, our measurements can be conservatively considered as upper limits, particularly for higher redshift targets among our sample ($z>0.1$).
To test the reliability of the size measurements, we apply the $\sigma_{\mathrm{[OIII]}}$-based method to various mock data. For this test, we set the redshift z=0.1 as a representative case and restrict $V_{max}$ from 750 to 2000 km s-1 to focus on strong outflows (see Section IV.2). In Figure 13, we present the measured outflow size as a function of the input (true) size. As expected, the outflow size is overestimated by more than a factor of two when the true outflow size is smaller than the FWHM of the seeing ($\sim$1.52 kpc at z = 0.1). On the other hand, when the seeing is smaller than the true size, the measured size is generally consistent with the input size with slight underestimations (<< 0.3 dex). Therefore, the $\sigma_{\mathrm{[OIII]}}$-based method provides a good size estimation within 0.3 dex, or at least can be used as a reliable upper limit, depending on the relative seeing size.
We emphasize that although the $\sigma_{\mathrm{[OIII]}}$-based method overestimates the sizes of compact outflows, the main conclusion of Kim et al. (2023), i.e., lack of global feedback, remains valid. Since the true size can be more compact than the measured one, the upper limit of the size still provides strong constraints on the spatial scales of AGN feedback.